# thanks: Corresponding author

### v1

@@ I Introduction
In recent years, the emergence of deep learning techniques has opened new avenues for efficient flow field predictions. A growing body of research has integrated deep learning methods with fluid flow simulation, yielding promising results. Notably, neural network architectures such as U-net Guo, Li, and Iorio (2016); Sekar et al. (2019); Ribeiro et al. (2020); Chen et al. (2021); Cai et al. (2022), generative adversarial networks (GANs) Goodfellow et al. (2020); Chen, Chiu, and Fuge (2020); Haizhou et al. (2022), and physics-informed neural networks (PINNs) Raissi, Perdikaris, and Karniadakis (2019); Cai et al. (2021); Chen, Liu, and Sun (2021); Zhang et al. (2023) have shown considerable potential in tasks such as flow field reconstruction. These approaches demonstrate the power of deep learning to capture the complex dynamics of fluid flow, signaling a shift in the future of CFD research. Guo et al.Guo, Li, and Iorio (2016) proposed a convolutional neural network (CNN)-based surrogate model for real-time prediction of steady laminar flow fields, achieving significantly faster velocity field estimation compared to traditional CFD solvers while maintaining low error rates. Wu et al.Haizhou et al. (2022) proposed a data-augmented Generative Adversarial Network (daGAN) for rapid and accurate airfoil flow field prediction, demonstrating strong generalization capabilities with sparse training data through pre-training and fine-tuning modules. Raissi et al.Raissi, Perdikaris, and Karniadakis (2019) introduced physics-informed neural networks , a deep learning framework that integrates physical laws governed by nonlinear partial differential equations to solve forward and inverse problems, enabling data-efficient spatio-temporal function approximation and scientific discovery even in small data regimes.
@@
Efficient and accurate prediction of high Reynolds number airflow around airfoils is crucial for advancements in aerodynamics research and the aviation industry, as it directly impacts airfoil optimization, optimal flow control, and the design of more efficient aerodynamic systems. However, due to the complexity of the flow—characterized by turbulence and flow separation—achieving both efficiency and accuracy in these predictions remains a significant challenge.
This study proposes a novel approach, Adaptive Diffusion Transformers for Airfoil Flow Simulation (AeroDiT), which utilize diffusion transformers to predict high Reynolds number airfoil flow by learning the data generated from RANS simulations. The primary objective is to improve the efficiency of airflow predictions using a generative surrogate model without compromising accuracy. The key contributions of this work can be summarized as follows.
First application of DiTs in aerodynamic flow prediction: This study pioneers the application of Scalable Diffusion Models with Transformers (DiTs) as surrogate models for flow field prediction, effectively addressing the prohibitive computational costs associated with conventional CFD methods.
Experimentally-validated accurate real-time prediction capability: The results demonstrate that the model achieves low relative L2 prediction errors in the experiment. Furthermore, the model enables real-time predictions in seconds, offering computational efficiency while accurately capturing complex flow characteristics around airfoils.
@@ III Methodology
Figure 1: The framework and learning process of the proposed AeroDiT. (a) The AeroDiT framework consists of a condition encoder, a function space decoder, and a transformer-based diffusion module. The encoder and decoder are asymmetrically trained. (b) Training process of the DiT model, illustrating the data flow through the AeroDiT architecture. (c) Inference process of the final model, where the desired output is generated from noise given condition inputs.
This study employs Diffusion Transformers to predict the airfoil flow field at high Reynolds numbers. The framework of DiTs is shown in Figure 1. By leveraging the generative nature of diffusion models, DiTs serve as a surrogate model that learns the statistical distribution of flow field data. Once trained, the model can predict velocity and pressure fields based on input airfoil geometry and flow conditions.
@@ IV.1.2 Configuration
Figure 3: Visualization of input data (Airfoil geometry,initial horizontal Velocity, initial parallel velocity), prediction (Pressure Field, velocity X, velocity Y) and the corresponding ground truth.
@@ IV.2 Results and Discussion
Table 1 summarizes the prediction results of the model using different DiT backbones. In experiments conducted on the annotated dataset, the best predicted $L_{2}$ errors were consistently below 0.105, 0.025, and 0.055, respectively, across all test cases.
Figure 3 presents the performance of the DiT-L/2 model trained on the Regular dataset for three different airfoil configurations, corresponding to Reynolds numbers of $2.04\times 10^{6}$, $1.06\times 10^{7}$, and $2.00\times 10^{7}$, representing the minimum, median, and maximum Reynolds numbers in the test set. Additionally, the differences between the predicted and ground-truth values are visualized. The discrepancies between predictions and ground truth are minimal and nearly indistinguishable to the naked eye. Furthermore, the model’s performance remains consistently reliable across both high and low Reynolds numbers, with negligible variation.
@@ IV.3 Case Study
Figure 7: Normalized flow field of AH 63-K-127/24 airfoil at $Re=9.11\times 10^{6}$
Figure 8: Comparison of surface pressure and velocities of Drela AG09 airfoil at $Re=9.11\times 10^{6}$
In test case 2, the airfoil is changed to AH 63-K-127/24 and the Reynolds number is increased to $9.11\times 10^{6}$. The normalized flow field of AeroDiTs and CFD in this case is given in Figure 7. Counterintuitively, the AeroDiTs predict better in this case compared to test case 1. The absolute errors of predictions are almost less than 0.075, 0.020 and 0.045 for pressure and velocity fields. Furthermore, Figure 8 gives the results of surface pressure and velocities. It can be seen that the overall curve match is also much better. One possible reason is that the training dataset contains more values around $Re=4.04\times 10^{9}$, making the model predictions more consistent with origin CFD results.
Figure 9: Normalized flow field of EPPLER 59 airfoil at $Re=1.21\times 10^{7}$.
Figure 10: Comparison of surface pressure and velocities of EPPLER 59 airfoil at $Re=1.2\times 10^{7}$.
Finally, the airfoil is selected to be EPPLER 59 and the Reynolds number is increased to $Re=1.21\times 10^{7}$. The comparison results of flow field and airfoil surface parameters are given in Figure 9 and Figure 10, respectively. From the prediction results, especially the airfoil surface parameters, the model under this test case does not perform well enough. There may be two reasons for this result. On the one hand, there are not enough cases of high Reynolds numbers in the training dataset. On the other hand, the flow field characteristics change drastically during the change of Reynolds number from the order of millions to the order of ten million. It may be possible to deal with this problem by adding more high Reynolds number data to the training dataset or by introducing physical equations to the neural networks.
@@ VI Conclusion
These findings highlight AeroDiT’s potential to bridge the gap between computational efficiency and predictive accuracy in CFD simulations. Future work should focus on enriching the training dataset with higher Reynolds number scenarios and incorporating physical constraints to improve model performance in complex turbulence regimes. Such integration would represent a critical step in advancing the capabilities of machine learning models for aerodynamics and other engineering applications.

### v2

@@ I Introduction
In recent years, the emergence of deep learning techniques has opened new avenues for efficient flow field predictions. A growing body of research has integrated deep learning methods with fluid flow simulation, yielding promising results. Notably, neural network architectures such as U-net Guo, Li, and Iorio (2016); Sekar et al. (2019); Ribeiro et al. (2020); Chen et al. (2021); Cai et al. (2022), generative adversarial networks (GANs) Goodfellow et al. (2020); Chen, Chiu, and Fuge (2020); Haizhou et al. (2022), and physics-informed neural networks (PINNs) Raissi, Perdikaris, and Karniadakis (2019); Cai et al. (2021); Chen, Liu, and Sun (2021); Zhang et al. (2023) have shown considerable potential in tasks such as flow field reconstruction. These approaches demonstrate the power of deep learning to capture the complex dynamics of fluid flow, signaling a shift in the future of CFD research. Guo et al.Guo, Li, and Iorio (2016) proposed a convolutional neural network (CNN)-based surrogate model for real-time prediction of steady laminar flow fields, achieving significantly faster velocity field estimation compared to traditional CFD solvers while maintaining low error rates. Raissi et al.Raissi, Perdikaris, and Karniadakis (2019) introduced physics-informed neural networks , a deep learning framework that integrates physical laws governed by nonlinear partial differential equations to solve forward and inverse problems, enabling data-efficient spatio-temporal function approximation and scientific discovery even in small data regimes.
@@
Efficient and accurate prediction of high Reynolds number airflow around airfoils is crucial for advancements in aerodynamics research and the aviation industry, as it directly impacts airfoil optimization, optimal flow control, and the design of more efficient aerodynamic systems. The application of deep learning in aerodynamics has evolved progressively from foundational flow field prediction to efficient optimization design. Early research (Sekar, 2019) established a speed foundation for replacing traditional CFD through a hybrid CNN-MLP architecture (where CNNs extract geometric features and MLPs map flow parameters), enabling real-time flow field prediction, though limited to steady laminar flows. To address data dependency issues, Wu (2021) introduced a Generative Adversarial Network (daGAN), leveraging a two-phase training mechanism (cGAN pre-training followed by dual-generator fine-tuning) to overcome sparse data constraints and enhance generalization under small-sample conditions. Li (2020) developed Deep Learning-based Geometric Filtering (DLGF), utilizing autoencoder-based compression of the design space to reduce variable dimensionality from hundreds to tens, thereby circumventing the curse of dimensionality and accelerating optimization workflows. Within this framework, Wang (2022) integrated a CNN-based Reduced Order Model (ROM) with a fusion strategy of pooling dimensionality reduction and radial basis functions to handle nonlinear flow optimization under large disturbances. Concurrently, Li (2021) addressed low-Reynolds-number separation bubbles by integrating the XFOIL solver with a transition-sensitive turbulence model, achieving efficient airfoil design for UAVs through tailored modal decomposition.
While these works mentioned above collectively aim to replace high-cost simulation iterations via a paradigm integrating physical constraints with data-driven intelligence, they exhibit limitations: conventional generative models (e.g., GANs, DDPMs) inadequately capture complex separated flows in high-Reynolds-number turbulence; U-Net-based diffusion models suffer from inefficient iterative sampling that precludes real-time application; and purely data-driven approaches show weak extrapolation capability under extreme Reynolds numbers while lacking embedded physical constraints, leading to localized errors near critical airfoil surfaces. However, due to the complexity of the flow—characterized by turbulence and flow separation—achieving both efficiency and accuracy in these predictions remains a significant challenge.
This study introduces a novel methodology for aerodynamic flow-field simulation—Adaptive Diffusion Transformers for Airfoil Flow Simulation (AeroDiT)—which employs diffusion transformers to learn from high-fidelity Reynolds-Averaged Navier–Stokes (RANS) simulation data. By treating the complex flow-field distributions as high-dimensional probability spaces and leveraging a generative approach, AeroDiT aims to dramatically reduce computational costs while preserving predictive accuracy at high Reynolds numbers. Compared to conventional computational fluid dynamics (CFD) techniques, AeroDiT’s innovations extend both the theoretical and practical frontiers of aerodynamic simulation, yielding the following key contributions:
1. Pioneering Application of Diffusion Transformers in Aerodynamic Flow Prediction. This work represents the first utilization of Scalable Diffusion Models with Transformers (DiTs) as surrogate models for predicting aerodynamic flow fields. Traditional CFD approaches to high-Reynolds-number turbulence require fine spatial discretization and iterative solvers, resulting in prohibitive computational overhead. AeroDiT reframes the flow-field as a high-dimensional latent distribution and embeds multiscale turbulent structures into a diffusion-based generative framework. The transformer’s self-attention mechanism captures interactions across all spatial scales, enabling efficient reconstruction of the entire flow field from learned statistical representations. From a theoretical standpoint, this extends the application paradigm of generative models within physical sciences and demonstrates a data-driven route to approximately solve partial differential equations (PDEs), thereby circumventing the heavy numerical discretization typical of classical methods.
2. Experimentally Validated, High-Fidelity, Real-Time Prediction Capability. During validation, AeroDiT exhibits exceptional accuracy and real-time performance. Across a variety of angles of attack, freestream velocities, and representative airfoil geometries, the model consistently achieves relative L2 prediction errors on velocity fields, pressure distributions, and turbulence metrics (such as vorticity) on the order of only a few percent. Simultaneously, the generative inference process can reconstruct a complete flow field in a matter of seconds—many orders of magnitude faster than RANS-based CFD. Consequently, AeroDiT holds immense promise for applications requiring massive parametric sweeps, real-time control, or rapid design iteration. It can accelerate the design optimization cycle in aerospace engineering and furnish reliable surrogate models for data-driven, multidisciplinary design optimization.
3. Integration of Physical Priors to Ensure Consistency with Governing Equations. Beyond empirical accuracy and speed, AeroDiT’s core innovation lies in blending generative modeling with physics-informed constraints. By embedding RANS-derived physical priors during training, the network continually corrects deviations from conservation laws, preventing unbounded extrapolation errors in out-of-distribution regimes. Such a methodology paves the way for extending generative deep learning to more complex flow problems (e.g., multiphase flows, combustion, three-dimensional wing arrays), offering a general framework for achieving a balance between physical fidelity and statistical expressivity.
@@ III Methodology
Figure 1: The framework and learning process of the proposed AeroDiT (Normalized flow field of EPPLER 59 airfoil).
This study employs Diffusion Transformers to predict the airfoil flow field at high Reynolds numbers. The framework of Diffusion Transformers (DiTs) is illustrated in Figure 1. By leveraging the generative capabilities of diffusion models, DiTs function as a surrogate model that effectively learns the underlying statistical distribution of flow field data. The model is designed to predict the flow characteristics, such as velocity components and pressure fields, based on input features including the geometry of the airfoil and various flow conditions. The overall architecture consists of three primary components: the Condition Encoder, the Diffusion Transformer (DiT) block, and the Function Space Decoder. The Condition Encoder processes input data such as the airfoil geometry and flow velocities to generate an initial latent representation. This encoded information is then fed into the DiT block, where noise is added, and a series of linear transformations, reshaping, layer normalization, and attention mechanisms take place to effectively model the underlying data distribution. The DiT block utilizes multi-head self-attention, pointwise feed-forward networks, and residual connections to capture the complex relationships between the variables in the latent space. Finally, the Function Space Decoder decodes the learned representation into meaningful output data, such as the predicted pressure field and velocity components (both horizontal and parallel), which are visualized in the final output.
Once trained, DiTs can be used as a predictive tool, providing accurate approximations of the flow field for different aerodynamic configurations, without the need for traditional computational fluid dynamics (CFD) simulations. This offers a computationally efficient and scalable alternative for real-time flow field predictions based on airfoil geometry and specified flow conditions. The geometric shape of the wing is shown in Figure 2, with blue for the flow field region and red for the wing shape.
@@ III.2 Learning Target
Figure 2: Wing geometry. (The red part shows the wing and the blue part shows the flow field)
@@ IV.2 Results and Discussion
Table 1 summarizes the prediction results of the model using different DiT backbones. In experiments conducted on the annotated dataset, the best predicted $L_{2}$ errors were consistently below 0.105, 0.025, and 0.055, respectively, across all test cases. The “$\pm$” values indicate the relative error percentage with respect to the mean $L_{2}$ error, providing a measure of variability across different test samples.
@@ IV.3 Case Study
The prediction errors on the airfoil surface are primarily concentrated in the pressure and transverse velocity fields, which are likely related to flow separation and boundary layer characteristics. In high Reynolds number flows, the turbulent structures and separation phenomena in the boundary layer significantly influence the flow field. AeroDiT may struggle with these complex flow features. From a physical perspective, these errors reflect the model’s limitations in capturing the fine details of the flow.
Moreover, AeroDiT achieves rapid predictions through its diffusion transformer architecture, significantly enhancing computational efficiency. However, this efficiency gain comes at the cost of some accuracy loss. While the model excels in predicting longitudinal velocity, errors persist in the pressure and transverse velocity predictions. This underscores the need to balance computational efficiency and prediction accuracy when designing deep learning models. For practical engineering applications, this trade-off should be adjusted according to specific needs. For instance, efficiency can be prioritized in the preliminary design phase, while higher accuracy is required in the optimization phase.
@@ VI Conclusion
These findings highlight AeroDiT’s potential to bridge the gap between computational efficiency and predictive accuracy in CFD simulations. Looking ahead, the continued development and refinement of AeroDiT hold the promise of even greater advancements. By enriching the training dataset with higher Reynolds number scenarios and integrating physical constraints, we can further enhance the model’s performance in complex turbulence regimes. This not only paves the way for more accurate simulations but also opens new avenues for exploring the intricate dynamics of airflow around airfoils.
As we stand on the cusp of a new era in computational fluid dynamics, AeroDiT stands as a testament to the transformative power of machine learning. Its potential applications in various fields, from aerospace engineering to renewable energy, are vast and far-reaching. By bridging the gap between computational efficiency and predictive accuracy, AeroDiT heralds a future where real-time aerodynamic simulations are not just a possibility but a reality.
Appendix A ref figures
Figure 7: Distribution of pressure and velocity fields around the airfoil from low to high Reynolds numbers for the same airfoil : $Re_{1}=9.11\times 10^{6}$, $Re_{2}=1.06\times 10^{7}$, $Re_{3}=2.00\times 10^{7}$
Figure 8: Distribution of pressure and velocity fields around different airfoils at the same high Reynolds number : $Re=1.21\times 10^{7}$.